Let f(x) = -x^3 + 3x^2 - 3x + 1, and g(x) be f(x) divided by 1 - x; solve for
g(x) if 1 -x is a factor or f(x).
a. g(x) = x^4 - 4x^3 + 6x^2 - 4x + 1
b. g(x) = x^3 - 3x^2 + 3x -1
c. g(x) = -x^2 + 2x - 1
d. g(x) = x^2 - 2x + 1

Answer by Edwin McCravy(20054)

Let f(x) = -x^3 + 3x^2 - 3x + 1, and g(x) be f(x) divided by 1 - x; solve for
g(x) if 1 -x is a factor or f(x).
a. g(x) = x^4 - 4x^3 + 6x^2 - 4x + 1
b. g(x) = x^3 - 3x^2 + 3x -1
c. g(x) = -x^2 + 2x - 1
d. g(x) = x^2 - 2x + 1


Write 1 - x as -x + 1 and divide by 
long division to find g(x):

              x² - 2x + 1
      ------------------- 
-x + 1)-x³ + 3x² - 3x + 1
       -x³ +  x²
       ---------
             2x² - 3x
             2x² - 2x
             --------
                   -x + 1
                   -x + 1
                   ------
                        0

So g(x) = x² - 2x + 1, which is 
choice d.

Edwin